OpenStudy (anonymous):

HELP PLEASE!! MEDAL & FAN!!! solve for x: A.)x = 2 B.)x = 6 C.)x = 41 D.)x = 43

OpenStudy (anonymous):

\[\frac{ 1-x }{ 2 }= \frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

help please: @ikatouni @Jadeishere @undeadknight26

OpenStudy (epoweritheta):

(1-x)/2=7/2 multiply both sides by 2 1-x=7 x=1-7=-6 none of the options given

OpenStudy (anonymous):

B.

OpenStudy (epoweritheta):

ikatouni B doesnt satisfy the equation

OpenStudy (anonymous):

the answer is 6 but it is a negative six I was going to right that

OpenStudy (anonymous):

hey, its ok... dont be arguing!! lolz

OpenStudy (anonymous):

gurl can u help me figure out which ones right?? @destinistmyers

OpenStudy (epoweritheta):

faylynn believe me : (

OpenStudy (anonymous):

@faylynn hey its -6 but I was going to write that until he interepted me

OpenStudy (anonymous):

ok... well thats still not one of the chooses tho, and idk whats the answer, would it be b??

OpenStudy (anonymous):

yes you are correct I just review the answer and u got it reat job :) girl power !

OpenStudy (epoweritheta):

even if -6 is not in option that is the answer

OpenStudy (anonymous):

thanks, is it right guys??

OpenStudy (anonymous):

@welly don't lie u got no show work you just want a medal and ur fake

OpenStudy (epoweritheta):

: (

OpenStudy (anonymous):

ok.. i get it already.. i am taking a flutter test, i cant just put -6 as the answer on a multiple choice question! and man u do that to everyone??

OpenStudy (anonymous):

@Monster1234 can u help??

OpenStudy (epoweritheta):

I am ePOWERitheta the super power , i always give the correrct answer lol : D

OpenStudy (anonymous):

OK.. but its still not one of the answers

OpenStudy (epoweritheta):

then the question is wrong , use wolfram alpha to solve

OpenStudy (anonymous):

omg im so sorry the answer is 43

OpenStudy (anonymous):

wtf?? wtf is that??

OpenStudy (anonymous):

r u messin silly goose??? @ikatouni

OpenStudy (anonymous):

no im serious

OpenStudy (anonymous):

6

OpenStudy (anonymous):

i think its 6

OpenStudy (anonymous):

@epoweritheta if u use wolframa u get -5

OpenStudy (anonymous):

@ikatouni can u show ur work please??

OpenStudy (epoweritheta):

everybody forgot their maths , answer is -6

OpenStudy (epoweritheta):

he cant show , because he just bluffed

OpenStudy (epoweritheta):

good night friends

OpenStudy (anonymous):

OMG..... how many times do i to tell u..... its a fluttering multiple choice question, if u got -6 it obviously wrong! @epoweritheta

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

43 is my final go by wat u want

OpenStudy (anonymous):

@Whitemonsterbunny17 @iPwnBunnies hey can u help.. PLEASE??

OpenStudy (anonymous):

@brandon1290

OpenStudy (anonymous):

@random231

OpenStudy (anonymous):

@geekfromthefutur

undeadknight26 (undeadknight26):

I am sorry but i got -6 also...I even tried it in a equation calculator... http://www.wolframalpha.com/input/?i=%28%281-x%29%2F2%29+%3D+7%2F2

OpenStudy (epoweritheta):

the options are wrong . -6 is correct

OpenStudy (anonymous):

ok.. i will go yell @ my teacher then!! thanks!!

OpenStudy (anonymous):

@faylynn I already told u it was 6 but I noticed it had to be negative and I was going to tell u until he interepted me

OpenStudy (anonymous):

Thanks ALL of u!! u guys r lifesavers!!! i know i said i was going to give a medal... but i not 'cause u ALL did great so i just going to fan!!!!! thanks SO much!!

OpenStudy (anonymous):

hey, @ikatouni if tag u in another math question, do u think u can help me?? i will give a medal!!

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok thanks SO much!!